two cars are traveling in the sam direction. one car is going 45 mph and the other car is going 60 mph, the first car left two hours ahead of the second car, explain how you could solve an equation to find out how long it will take the second car to catch up to the third
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

Car I speed= 45 mph , starts at
Car II speed= 60 mph starts at

Time difference between the starting of the two= 2.00 hours

Time taken by Car I = X / 45
Time taken by Car II = X / 60

X / 45 - X / 60 = 2

LCD - 90
Multiply equation by 90 90

2 x - 1.5 x = 180
0.5 x = 180
/ 0.5
x= 360 miles
Car II speed= 60 mph
Distance = 360 miles
Time taken = 360 / 60
Time taken = 6 hour

Answer by richwmiller(17219)

Gaining speed 60 -45= 15 mph
Head start 45*2 =90 miles
Time to catch up 90/15=6 hours
Algebraic solution
45(t+2)=60t
45t+90=60t
90=60t-45t
90=15t
t=6 hours
Distance to meeting 6*60=360 miles
code drtch1